A study conducted on​ people's main source of news was based on 250 respondents between the ages of 36 to 50 years and 250 respondents who were over age 50. Of the 250 respondents who were between the ages of 36 and​ 50, 81 got their news primarily from newspapers. Of the 250 respondents who were over age​ 50, 157 got their news primarily from newspapers. Complete parts​ (a) through​ (d).
36-50 Over 50 Total
Yes 81 157 238
No 169 93 262
Total 250 250 500
b. What is the probability that an individual respondent is over 50 years​ old?



Answer by math_tutor2020(3835)


	36-50	Over 50	Total
Yes	81	157	238
No	169	93	262
Total	250	250	500


There are 250 people over age 50. Look at the bottom of the "over 50" column to see this total. Let A = 250.

There are 500 people overall (bottom right corner). Let B = 500.

The probability of picking someone over 50 is A/B = 250/500 = 1/2.

Answer: 1/2